OpenStudy (anonymous):

Given that f(x) = 3x + 1 and g(x) = 4x+2 all over 3 solve for g(f(0))

OpenStudy (anonymous):

2 6 8 9

OpenStudy (btaylor):

f(0) = 3(0) + 1 = 1 g(1) = 4(1) + 2 = 6

OpenStudy (anonymous):

so the answer is 6 @BTaylor

OpenStudy (btaylor):

yep.

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

could you help with another one @BTaylor

OpenStudy (anonymous):

The function b(x) = 2x - 5 determines how many bags of dog food need to be purchased for an animal shelter, where x is the number of dogs at the shelter. The shelter manager uses m(b(x)) to find the amount of money to bring for the dog food purchase. The function m(x) = 3x + 1. Solve for how much money to bring when there are 10 dogs in the shelter. 15 31 46 52

OpenStudy (btaylor):

plug 10 into b(x), then plug that result [b(10)] into m(x)

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

b(x)=15

OpenStudy (anonymous):

idk what to do next

OpenStudy (anonymous):

@tester97 help please

OpenStudy (btaylor):

plug 15 into m(x)

OpenStudy (btaylor):

you get 46.